HUBBLE SPACE TELESCOPE DAILY REPORT #2858 PERIOD COVERED: 0000Z (UTC) 04/24/01 - 0000Z (UTC) 04/25/01 Daily Status Report as of 115/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 8827 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed nominally. 1.2 Completed Two Sets of WF/PC-2 8811 (Cycle 9 Standard Darks) The WF/PC-2 was used to obtain dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. Over an extended period these data will also provide a monitor of radiation damage to the CCDs. The proposal completed with no reported problems. 1.3 Completed Two Sets of STIS/CCD 8864 (CCD Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks for the CCD. The proposal completed nominally. 1.4 Completed WF/PC-2 8581 (A Search For Low-Mass Companions To Ultracool Dwarfs) The WF/PC-2 was used to search for very low-mass {VLM} companions to a complete sample of 120 late-M and L dwarfs, drawn mainly from the 2MASS and SDSS surveys. The primary goal is to determine the multiplicity of M < 0.1 M_odot dwarfs. In particular, we aim to identify binary systems suitable for long-term astrometric monitoring and mass measurement, and systems with cool, sub-1000K companions. The proposal completed with no reported problems. 1.5 Completed Three Sets of STIS/CCD 9248 (Probing the Large Scale Structure: Cosmic Shear Observations) The Space Telescope Imaging Spectrograph (CCD) was used to measure the distortion of light bundles from distant galaxies that probe the statistical properties of the intervening inhomogeneous {dark} matter distribution. Its tidal gravitational field distorts the observable image shapes thereby causing a coherent ellipticity pattern {Cosmic Shear}. The loss-of-lock described in 2.1 and HSTAR 8183 occurred during the second iteration of this observation. Otherwise, the observations completed nominally. 1.6 Completed STIS/CCD 8594 (H-Deficient Condensations In Pne -- A Key To Discrepancies In Abundance Determinations) The Space Telescope Imaging Spectrograph (CCD) was used to obtain deep long-slit spectra for the well-resolved H-deficient knots of A 30 and to search for such knots. The data will yield the spatially resolved temperature, density and ionization structure of the knots in Abell 30, and, for the first time, accurate ORL C and O abundances for them. The results will lead to a much better understanding of the physics of such knots and their effects on abundance determinations . There were no reported problems. 1.7 Completed Four Sets WF/PC-2 8816 (Cycle 9 UV Earthflats) The WF/PC-2 was used to obtain sequences of Earth streak flats to improve the quality of pipeline flat fields for the WFPC2 UV filter set and in order to monitor flat field stability. There were no reported problems. 1.8 Completed Eight Sets of WF/PC-2 9244 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The loss-of-lock described in 2.1 and HSTAR 8183 occurred during the second iteration of this observation. Otherwise, the observations completed with no other anomalous activity. 1.9 Completed FGS/1 9240 (Astrometric Mass Determination of an Extrasolar Planet Candidate) Fine Guidance Sensor-1R was used to perform astrometric observations that will allow a quick measure of the sky-plane motion of 55 Cancri, a G star with an Msin{i} = 0.9 MJ radial velocity companion with a 14 day period. By using 7 pairs of visits, where each visit within a pair has the same parallax factor, we will be able to accurately determine the stars's proper motion. The reflex motion of the star, due to its orbit about the star-companion barycenter, will manifest itself as a systematic residual to the measured proper motion if the companion is sufficiently massive. The loss-of-lock described in 2.1 and HSTAR 8183 occurred during this observation. There were no other reported problems. 1.10 Completed STIS/CCD 8865 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.11 Completed WF/PC-2 8682 (A Snapshot Study of 0bservational Cosmology) The WF/PC-2 was used to examine the observational constraints on the cosmic star formation history that is currently among the most active fields in observational cosmology. The most widely used tracer of the co-moving volume-averaged star formation rate {SFR} is the UV luminosity density, which early results found to peak at z~1- 2. The apparent identification of the primary epoch of metal production and star formation in the Universe led to intense theoretical and observational interest. Nevertheless, and remarkably for such a fundamental observation, little is known about the history of star formation in the Universe beyond its global average. There were no reported anomalies. 1.12 Completed WF/PC-2 8634 (Atmospheric Variability on Uranus and Neptune) The WF/PC-2 was used to obtain snapshot observations of Uranus and Neptune to monitor changes in their atmospheres on time scales of months to years. Uranus is rapidly approaching equinox in 2007, with another 4 degrees of latitude in the northern hemisphere becoming visible every year. Recent HST images during this unique epoch have revealed: {i} strongly wavelength-dependent latitudinal structure, {ii} the presence of numerous visible-wavelength cloud features in the northern hemisphere, and, {iii} in the near-IR, discrete features northward of 25degrees N that have the highest contrast ever seen for a Uranian cloud. Long-term ground- based observations show seasonal brightness changes whose origins are not well understood. Recent IR images of Neptune obtained using adaptive optics on the Keck Telescope indicate that a new "Bright Companion" type of feature has recently appeared in the southern hemisphere. Snapshot observations of these two dynamic planets can supply context in which to discern the nature of long-term changes in their latitudinal atmospheric bands and to track the appearance, movement, and disappearance of discrete albedo features. There were no reported problems. 1.13 Completed Three Sets of STIS/MA1/MA2 8843 (Cycle 9 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to perform the routine monitoring of the MAMA detector dark noise. The proposal completed nominally. 1.14 Completed WF/PC-2 8712 (The Wind Accretion Wake In A Detached Binary System) The WF/PC-2 was used to observe RW Hya {= HD 117970}, the only known detached binary system, where direct observational evidence for wind accretion onto a white dwarf is found. UV observations will enable the investigator to advance knowledge in a field which has for long had to remain a theoretical subject. RW Hya is an eclipsing system containing a mass-losing M-giant and a hot white dwarf on circular orbits {370 days}. The observation completed with no reported problems. 1.15 Completed STIS/CCD/MA1/MA2 8635 (A Test Of Pulsation And Diffusion Theory For Subluminous B Stars) The Space telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to confirm the recent discovery of radial and nonradial mode pulsations in nearly 20 sdB stars that makes it possible to use asteroseismology to probe the internal structure of these stars and discern their evolutionary status. This is needed for reasons as diverse as understanding the late stages of stellar evolution and the calibration of the observed ultraviolet upturn in giant elliptical galaxies as an age indicator. The observations completed with no anomalous activity. 1.16 Completed STIS/CCD/MA1 8662 (A Snapshot Survey of the Hot Interstellar Medium) The Space Telescope Imaging Spectrograph (CCD And MA1) was used to obtain snapshot STIS echelle observations of key tracers of hot interstellar gas {C IV, N IV, and Si IV} for selected FUSE Team O VI survey targets with known UV fluxes. By taking advantage of the snapshot observing mode we will efficiently obtain a large number of spectra suitable for the study of the highly ionized hot component of the interstellar medium {ISM}. Our goals are to explore the physical conditions in and distribution of such gas, as well as to explore the nature of the interfaces between the hot ISM and the other interstellar gas phases. The observations completed with no reported problems. 1.17 Completed STIS/CCD/MA2 8595 (Does the D/H Ratio Vary in Local Interstellar Gas?) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to take measurements of the D/H ratio in the Local Interstellar Cloud {LIC} that are consistent with 1.5 +/- 0.1 X 10^-5, while for other clouds within ~ 100 pc, the measurement uncertainties are larger. Two target white dwarfs {WDs}, HZ 43 and GD 153, are chosen because EUVE data show low HI column densities, thereby reducing the probabilities of multiple velocity components and saturation of the DI line. The WDs have pure-H atmospheres, so contamination of the interstellar lines by photospheric metal lines is not an issue. The measurements completed with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 11 Successful: 11 After a successful acquisition, there was a loss-of-lock at 114/122553Z during an astrometry observation. Recovery was not completed until 114/124411Z. HSTAR 8183 was written. Proposals detailed in 1.5, 1.8 and 1.9 were affected. Scheduled Re-acquisitions: 4 Successful: 4 2.2 FHST Updates: Scheduled: 30 Successful: 30 2.3 Operations Notes: Using ROP SR-1A, the SSR EDAC error counter was cleared twice. The CCS engineering status buffer limits were adjusted five times as directed by ROP DF-18A. Per ROP RD-7A, ESTR-2 reconditioning was completed at 114/0917Z. At 114/122506Z, per HSTAR 8182, while guiding, the gyro #5 motor current spiked to 254.8 mAmps and then settled down to a toggle between 204 mAmps and 210 mAmps. This is the second current spike on this gyro in two days. This event occurred during the loss-of-lock event described in 2.1 above. The spike was only six mAmps from the 270 mAmp threshold which would have caused the gyro to be taken offline, resulting in a subsequent Zero Gyro safemode entry. Initial indications are that there was a decrease in the gyro duty cycle and a bias shift increase. The gyro #5 motor current operational yellow limits were again adjusted upward at 114/2258Z per an operations request. At a meeting held at 3:00 p.m., it was decided to perform data extractions on the original gyro #4 motor current, the origianl gyro#5 spin down, and motor currents for all prior gyro power ups. PCS and flight software personnel were given the action to modify the gyro #5 motor current limit test to reflect 299.6 mAmps. SSA transmitter #2 was turned on at 114/1739Z and turned off at 114/1748Z, using ROP IC-2. TTRs were generated for required re-transmits during at NSSC-1 uplink at 115/0416Z and during a 486 uplink at 115/0735Z. After the first occurrence, using ROP NS-5, a NSSC-1 error was reset at 115/0424Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW